# Understanding Self-Admitted Technical Debt in Test Code: An Empirical Study

### v1

@@ 6.3. Implications
          RQ2 clarified that methods containing Test SATD have more lines of code (LOC), annotations, and code smells than methods without Test SATD. From the perspectives of LOC and annotations, this suggests that Test SATD tend to reside in more complex test code. However, we also found that Test SATD is not associated with test smells (i.e., they are independent). This implies that test smell detection tools are not identifying most of the test issues, which is a gap to be fill by researchers and tool developers. Compared with code smells, test smells are a relatively recent concept and are still under development. Therefore, we recommend that researchers and tool developers should endeavor to identify new types of test smells, referring to Test SATD in practice.
@@ 7. Threats to validity
          The imbalanced distribution across categories poses validity concerns. The most infrequent category, “Failure,” contains only 15 instances, which may not adequately represent the full spectrum of this construct. While this imbalance problem is common in software engineering research (Seiffert et al., 2014; Tantithamthavorn et al., 2020), this imbalance directly affects the validity of our performance measurements: during 10-fold cross-validation, several test sets contained only a single “Failure” instance, potentially leading to unstable performance estimates that may not accurately reflect the model’s true ability to identify this category. While using median performance metrics partially addresses this threat, we cannot claim equal measurement validity across all categories.

### v2

@@ 6.3. Implications
          RQ2 clarified that methods containing Test SATD have more lines of code (LOC), annotations, complexity, and code smells than methods without Test SATD. From the perspectives of LOC, annotations, and complexity, this suggests that Test SATD tends to reside in larger and more complex test code. However, we also found that Test SATD is not associated with test smells (i.e., they are independent). This implies that test smell detection tools are not identifying most of the test issues, which is a gap to be filled by researchers and tool developers. Compared with code smells, test smells are a relatively recent concept and are still under development. Therefore, we recommend that researchers and tool developers should endeavor to identify new types of test smells, referring to Test SATD in practice.
@@ 7. Threats to validity
          The imbalanced distribution across categories poses validity concerns. The most infrequent category, “Failures,” contains only 15 instances, which may not adequately represent the full spectrum of this construct. While this imbalance problem is common in software engineering research (Seiffert et al., 2014; Tantithamthavorn et al., 2020), this imbalance directly affects the validity of our performance measurements: during 10-fold cross-validation, several test sets contained only a single “Failures” instance, potentially leading to unstable performance estimates that may not accurately reflect the model’s true ability to identify this category. While using median performance metrics partially addresses this threat, we cannot claim equal measurement validity across all categories.